OpenStudy (anonymous):

Trevor is tiling his bathroom floor, which has an area that is represented as 117r4 square inches. Each tile has an area of square root of the quantity 9 r to the thirteenth power . The total number of tiles used can be represented by the expression below. one hundred seventeen r to the fourth power, all over the square root of the quantity nine r to the thirteenth power Simplify the expression for the total number of tiles used. Show your work. Idk how to do it, help ?

OpenStudy (anonymous):

I really need help soon if possible :/

OpenStudy (anonymous):

Here is the equation:

OpenStudy (anonymous):

I would start by getting rid of the square root sign So square it all, let me know what you get :)

OpenStudy (anonymous):

Square the bottom? or the top? or both?

OpenStudy (anonymous):

what you do to one you must do to the other so yep, to both

OpenStudy (anonymous):

Would I square the r's?

OpenStudy (anonymous):

Would it be 20,709r^4 / sqrt 81r^13?

OpenStudy (anonymous):

@sarahusher ?

OpenStudy (anonymous):

Yes you would square the r's, so you should get \[\frac{ (117r ^{4})^{2} }{ (\sqrt{9r ^{13}})^{2} }=\frac{ 13689r ^{8} }{ 9r ^{13} }\]

OpenStudy (anonymous):

Okay what do I do now?

OpenStudy (anonymous):

so ignoring the numbers for now, if you have 8 'r's' on the top, and 13 'r's' on the bottom, how can you simplify this?

OpenStudy (anonymous):

You subtract them?

OpenStudy (anonymous):

^^ by the way, when you square a 'square root' the square root goes away, but the number inside stays the same

OpenStudy (anonymous):

Okay got it. So you subract the r's so it's r^13 -r^8, so it's r^5?

OpenStudy (anonymous):

yes! so as the 13 is on the bottom (ie it's the bigger one) you will still have 5 lots of 'r' left on the bottom so now you have\[\frac{ 13689 }{ 9r ^{5} }\] now can you simplify those numbers?

OpenStudy (anonymous):

I'm not sure how to

OpenStudy (anonymous):

okay i'll simplify it for you a bit 13689 = 9*1521 So this equation above ^^ can be re written as\[\frac{ 13689 }{ 9r ^{5} }=\frac{9*1521 }{ 9*r ^{5} }\] Can you see any common factor that you can cancel out?

OpenStudy (anonymous):

The 9?

OpenStudy (anonymous):

yep! so what would you get after doing that?

OpenStudy (anonymous):

1521/r^5 ?

OpenStudy (anonymous):

Great! That's it! Good work!

OpenStudy (anonymous):

Thank's so much! You were great help ! :)

OpenStudy (anonymous):

You're very welcome!